OpenStudy (anonymous):

A disk has a constant angular acceleration of 4.00 rad/s^2 about an axis perpendicular to the disk at its center. Find the radius of a point on the disk, where 0.500s after the disk begins to rotate, the magnitude of the total acceleration (centripetal plus tangential) equals that of the acceleration due to gravity. I still don't understand.

OpenStudy (experimentx):

\( a_t = r \alpha \sin\theta \) , if \( \theta =\pi/2\) , \( a = r \alpha \) <-- tangential acceleration at certain point at distance 'r' \( a_c = v^2/r\), and \( v = \alpha t\) if \( v = 0 \) at \( t = 0 \)

OpenStudy (experimentx):

magnitude of resultant acceleration at certain point = \( \sqrt{a_c^2 + a_t^2} \) and for the last part ... I am not really sure at all!!!

OpenStudy (anonymous):

thank you.